R. Bachimanchi, K. Baggett, G. Ciovati, H. Fanning, C. Hovater, G. Krafft, A. Kimber, K. Macha, Ju. Musson, J. Musson, T. Plawski, T. Powers, R. Rimmer, and C. Reece.
Agenda and Notes
1. Operations Status and Summary/Review (Bickley)
Low energy setup with reduced heating accomplished. Though overall month for ops but not so much SRF/RF related. Lots of difficulties with machine reproducibility, optics, etc., although some problems with separators. Seem to be moving now. 1L06 energy drain from bad control card had been limiting the current. Current limit possibly gone now.
Plan for module swaps complete.
2. Data Extraction LLRF (Julia Musson)
3. Some Comments on particulates (Geng)
4. Plasma Processing (Powers)
Leadership of quenches. Offline discussion.
C100 Refurbishment Project for the future?